OpenStudy (anonymous):

A honda civic moving at 100 km/h is 80km behind a Ford moving in the same direction at 60 km/h. How far does the Honda civic travel before it catches up with the ford?

OpenStudy (anonymous):

Let h be hours: \[100(h)=60(h)+80\]\[40(h)=80\]\[h=2\] Now you know that it will take two hours to catch up. Insert 2 for 100h and you get 200 miles. It will take 2 hours and 200 miles to catch up with the Ford.

OpenStudy (anonymous):

Thanks a lot. :)

OpenStudy (anonymous):

Given, velocities v1=100 km/h , v2=60 km/h and distance d=80 km v=d/t or t=d/v therefore, t= d/v1-v2 t=80 / (100-60) = 80 / 40 = 2hr d=v x t = 100 x 2 = 200 km So, honda civic travels 200km before it catches up with the ford and it takes 2hours since honda civic travels 100km/hr.